i don't understand how to find the slope and y intercept of 5x-3y=30??????
Answer by LarissaRichardson(34)

You need to start by putting it slope intercept form: y=mx+b
5x-3y=30 I suggest to keep the y value positive if possible.
In this case:
3y=5x-30
Now divide by 3 to get:
y=5\3x-10
m=slope
b=y intercept
Hence:
slope=5/3
y-intercept=-10